Suppose b > d +1/3 ,c+1<a – 4, and d + 5/8> a + 2. Order a, b, c, and d from greatest to least. Explain your reasoning.
1
Expert's answer
2020-11-30T09:34:30-0500

Let's first write down properties of order used here :

  1. "x>y, z>h" implies "x+z>y+h"
  2. "x>y, y>z" implies "x>z"

Now let's analyze every inequality we have :

  • "b>d+\frac{1}{3}" implies "b>d" , as "\frac{1}{3}>0" and thus "b > d+\frac{1}{3} >d"
  • "c+1<a-4" implies "c<a", as "c+1<a-4" implies "c<a-5<a"
  • "d+\frac{5}{8}>a+2" implies "d>a" , as "d+\frac{5}{8} > a+2" implies "d>a+1 \frac{3}{8}>a"

Thus we have "b>d, d>a, a>c" so we can conclude that "b>d>a>c" and the ordering from the greatest to the least is "b,d,a,c" .